A 12 kg block slides down the entire length of a 2.3 m long ramp with a coefficient of kinetic friction of 0.37. The work done by friction on the block is 94 J. What is the angle of inclination of the ramp?


1
Expert's answer
2021-07-13T13:34:21-0400
"W_{fr}=F_{fr}d=\mu_kmgdcos\theta,""\theta=cos^{-1}(\dfrac{W_{fr}}{\mu_kmgd}),""\theta=cos^{-1}(\dfrac{94\ J}{0.37\cdot12\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot2.3\ m})=20^{\circ}."